A sample of 49 observations is taken from a normal population with a standard deviation of 10. The sample mean is 55. 

Determine the 99% confidence interval for the population mean. (Round your answers to 2 decimal places.) 

Confidence interval for the population mean is _______ and _______ .


A research firm conducted a survey to determine the mean amount Americans spend on coffee during a week. They found the distribution of weekly spending followed the normal distribution with a population standard deviation of $5. A sample of 49 Americans revealed that X̅ = $20. 

a. What is the point estimate of the population mean?

b. Using the 95% level of confidence, determine the confidence interval for u. (Round your answers to 2 decimal places.) 
Answer #1

ElO- ) Solutions Given n=49 6=10 X=55o. gg.% confidence internal for the Construct mean population 6.5 = (x + Za ( 6 ) =(5s += (ss [ 2576 (H)) (5st 3.68) =(55–368 , 5573.68 ) = (51:32 , 58.68 ) lorentini confidence interval for the populatiun mean 51= ( 2011:96 ()) 49 = (20 + 1.4) = (18.6 , 21.4) te confidence interval for the population mean is (18.6) and 1214 |